# Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements

### v1

@@ B.1 Key Distance Ladder Measurements
Gaia EDR3 parallaxes to Milky Way Cepheids (corrected for zero-point offset)
Megamaser geometric distance to NGC 4258 [Reid2019]
Detached eclipsing binary distances to the LMC [Pietrzynski2019]
@@
The 2022 HST result [Riess2022] yielded km s-1 Mpc-1 (1.4% precision). The 2025 JWST result [Riess2025] yields km s-1 Mpc-1 (1.3% precision), confirming the HST photometry and ruling out crowding-induced biases as the source of the tension.
@@
The 2021 analysis [Freedman2021] yielded km s-1 Mpc-1, intermediate between Planck and SH0ES. This sparked debate about whether TRGB measurements genuinely favor lower or whether methodological differences explain the offset from SH0ES.
The 2024 JWST analysis [Freedman2024] found km s-1 Mpc-1 from Cepheids, now more consistent with SH0ES. The CCHP continues to analyze TRGB and JAGB calibrations from JWST data.
@@
The Lee et al. [Lee2024] JAGB analysis yields km s-1 Mpc-1, consistent with Planck and below other distance ladder measurements. This represents the most significant outlier within Category 1.
The JAGB method uses carbon stars on the Asymptotic Giant Branch, which exhibit a narrow luminosity distribution in the J-band. The discrepancy with Li et al. [Li2024b] () appears to originate from differences in:
@@
The measurement by Wagner et al. [wagner2025] (2025) reports km s-1 Mpc-1 using the local expansion parameter , based on the infall velocity relation . However, the choice of appears somewhat arbitrary; standard models of the local Hubble flow typically adopt . Since , adopting —a value intermediate between the theoretical expectation and the assumed value—would yield km s-1 Mpc-1, shifting the result toward the Planck value and closer to the Distance Ladder mean value. The systematic uncertainty in propagates directly into but this source of error was not included in the original analysis [wagner2025]. However, we do include this effect in our analysis and the data shown in Table 9.
@@
Zhang et al. (2024) [Zhang2024] measured the Hubble constant using galaxy pairwise peculiar velocities from the Cosmicflows-4 catalog, obtaining km s-1 Mpc-1 (statistical). However, as discussed in [Zhang2024], several systematic uncertainties are not fully accounted for in the reported error budget. The distance calibration of CF-4 galaxies introduces uncertainties of 2–3%, translating to 1.5–2.3 km s-1 Mpc-1 in . The halo mass matching between simulations and observations, which is crucial for their method, contributes an estimated systematic error of 1–2% (0.8–1.5 km s-1 Mpc-1). Additionally, the modeling of pairwise velocities in the nonlinear regime, where theoretical predictions are less certain, introduces uncertainties of 2–3% (1.5–2.3 km s-1 Mpc-1). The choice of the maximum separation scale ( Mpc) and the selection of the velocity threshold add further systematic uncertainties of order 1% (0.8 km s-1 Mpc-1). Cosmic variance, despite mitigation through multiple simulation realizations, remains a source of systematic error. Combining these contributions in quadrature yields a total systematic uncertainty of approximately 3–4 km s-1 Mpc-1, which is comparable to or larger than the quoted statistical uncertainty. Therefore, a more complete error budget would be km s-1 Mpc-1, or equivalently km s-1 Mpc-1 when statistical and systematic errors are combined in quadrature. This corrected level of uncertainty is used in our analysis.
B.2 Key One-Step Measurements
@@ B.2.1 TDCOSMO: Evolution of Strong Lensing Constraints
TDCOSMO IV (Birrer et al. 2020): [Birrer2020] Incorporating stellar kinematics to constrain the mass-sheet degeneracy (MSD) shifted the result to km s-1 Mpc-1—now consistent with Planck. This dramatic shift demonstrated that power-law profiles alone cannot break the MSD.
Recent results (2023–2025): Subsequent analyses using hierarchical modeling, additional lens systems, and improved kinematics yield in the range – km s-1 Mpc-1, depending on modeling assumptions [Shajib2023, Birrer2025].
@@
We exclude the original Pesce et al. measurement from our sample, as the more recent Barua et al. analysis uses the same underlying data while addressing the known systematic uncertainties of the earlier measurement
B.2.2 Kelly et al. (2023): SN Refsdal Time Delays
@@
Supernova Refsdal, the first multiply-imaged supernova [Kelly2015], provides unique advantages for time-delay cosmography: the supernova light curve allows unambiguous identification of multiple images, unlike quasar variability.
@@
The Kelly et al. [Kelly2023] analysis uses time delays from the “Einstein Cross” configuration plus the predicted reappearance of image “SX.” The result, km s-1 Mpc-1, is consistent with Planck.
Follow-up analyses with different cluster mass models yield consistent results: Li et al. [Li2024a] find km s-1 Mpc-1. The consistency across mass models provides confidence in the robustness of this constraint.
B.2.3 Piratova-Moreno et al.:FRBs
@@
Piratova-Moreno et al. (2025) [PiratovMoreno2025] employed Fast Radio Bursts (FRBs) as cosmological probes to measure the Hubble constant using a catalog of 98 localized FRBs. Using the Maximum Likelihood Estimate (MLE) method with confirmed FRBs, they obtained km s-1 Mpc-1 (statistical), while their mock FRB catalog yielded km s-1 Mpc-1. However, this method is subject to substantial systematic uncertainties that are not fully quantified in the reported error budget. The primary systematic error arises from the decomposition of the observed dispersion measure (DM) into its constituent components: the Milky Way interstellar medium and halo (), the intergalactic medium (), and the host galaxy (). The uncertainty in depends on the Galactic electron density model used (NE2001 vs. YMW16), with differences of 20% at low Galactic latitudes, contributing 1.5–2 km s-1 Mpc-1 to the uncertainty. The host galaxy contribution, , is highly uncertain and varies dramatically between FRBs—from 30 pc cm-3 for typical hosts to over 900 pc cm-3 for extreme cases like FRB 20190520B. Studies suggest uncertainties of 50–100 pc cm-3, which translate to systematic errors of 2–3 km s-1 Mpc-1 in . Additionally, the baryon fraction in the IGM () introduces further systematic uncertainty; recent estimates range from to , corresponding to 1–2 km s-1 Mpc-1 uncertainty in . The choice of DM– relation (linear vs. power-law) also significantly impacts results, with Piratova-Moreno et al. reporting values ranging from km s-1 Mpc-1 (linear) to km s-1 Mpc-1 (power-law), indicating substantial model-dependent systematic uncertainties of order 5–15 km s-1 Mpc-1. Combining these systematic contributions in quadrature yields a total systematic error of approximately 6–8 km s-1 Mpc-1 for the MLE result, dominating over the statistical uncertainty. A more realistic error budget would therefore be km s-1 Mpc-1, or equivalently km s-1 Mpc-1 when combined in quadrature, reflecting the current limitations of FRB-based cosmology. This is the value used in our analysis.
B.2.4 Vogl et al. (2024): Tailored EPM
@@
The Vogl et al. [Vogl2024] tailored Expanding Photosphere Method represents the most precise model-independent measurement supporting the high- solution outside the traditional distance ladder.
@@
B.2.5 Barua et al. (2025): Megamaser Cosmology Project
The Megamaser Cosmology Project measures geometric distances to galaxies with circumnuclear water maser disks. The Barua et al. [Barua2025] frequentist analysis yields km s-1 Mpc-1.
@@
B.2.6 Pogosian et al. (2024): Sound-Horizon-Free BAO
The Pogosian et al. [Pogosian2024] analysis demonstrates that can be constrained from BAO data without assuming the CMB Sound Horizon.
@@
Results: DESI BAO yields km s-1 Mpc-1; pre-DESI BAO yields km s-1 Mpc-1.
Significance: These measurements anchor the low- solution in Category 4. Their high precision (%) means they dominate the weighted mean of this category, contributing to the very low .
B.2.7 Gravitational Wave standard sirens
GW standard sirens provide constraints independent of electromagnetic distance ladders:
Bright sirens (GW170817): The binary neutron star merger with optical counterpart yields km s-1 Mpc-1 [Palmese2023], limited by the peculiar velocity of NGC 4993 and viewing angle degeneracy.
Dark sirens (GWTC-3): Statistical analysis of 47 events without counterparts yields km s-1 Mpc-1 [Abbott2021], consistent with Planck.
@@ B.2.7 Gravitational Wave standard sirens
B.2.8 Fast Radio Burst Measurements
@@
Results: Published values span – km s-1 Mpc-1. Maximum-likelihood approaches [PiratovMoreno2025] tend to favor lower values ( km s-1 Mpc-1).
As FRB samples grow and host DM distributions become better characterized, this probe may provide competitive constraints.
@@ C.1.1 Distance Ladder Correlations
SNe Ia samples: Hubble-flow SNe Ia are predominantly from Pantheon+ [Brout2022], with substantial overlap. standardization parameters (, , ) are often shared.
@@ C.2 Methodology for Correlation-Adjusted Estimates
where is the covariance matrix with elements .
For simplicity, we adopt a uniform correlation structure within each category:
@@ C.3.1 Distance Ladder (Category 1)
compared to km s-1 Mpc-1 under independence.
The central value shifts by only km s-1 Mpc-1 (within statistical uncertainty), but the uncertainty increases by %.
C.3.2 Combined One-Step (Categories 2+3+4)
@@
compared to km s-1 Mpc-1 under independence.
The central value shifts by km s-1 Mpc-1 , and the uncertainty increases by %.

### v2

@@ B.1 Key Distance Ladder Measurements
The 2022 HST result [150] yielded km s-1 Mpc-1 (1.4% precision). The 2025 JWST result [149] yields km s-1 Mpc-1 (1.3% precision), confirming the HST photometry and ruling out crowding-induced biases as the source of the tension.
@@
The 2021 analysis [63] yielded km s-1 Mpc-1, intermediate between Planck and SH0ES. This sparked debate about whether TRGB measurements genuinely favor lower or whether methodological differences explain the offset from SH0ES.
The 2024 JWST analysis [62] found km s-1 Mpc-1 from Cepheids, now more consistent with SH0ES. The CCHP continues to analyze TRGB and JAGB calibrations from JWST data.
@@
The Lee et al. [96] JAGB analysis yields km s-1 Mpc-1, consistent with Planck and below other Distance Ladder measurements. This represents the most significant outlier within Category 1.
The JAGB method uses carbon stars on the Asymptotic Giant Branch, which exhibit a narrow luminosity distribution in the J-band. The discrepancy with Li et al. [98] () appears to originate from differences in:
@@
The measurement by Wagner et al. [183] (2025) reports km s-1 Mpc-1 using the local expansion parameter , based on the infall velocity relation . However, the choice of appears somewhat arbitrary; standard models of the local Hubble flow typically adopt . Since , adopting —a value intermediate between the theoretical expectation and the assumed value—would yield km s-1 Mpc-1, shifting the result toward the Planck value and closer to the Distance Ladder mean value. The systematic uncertainty in propagates directly into but this source of error was not included in the original analysis [183]. However, we do include this effect in our analysis and the data shown in Table 10.
@@
Zhang et al. (2024) [193] measured the Hubble constant using galaxy pairwise peculiar velocities from the Cosmicflows-4 catalog, obtaining km s-1 Mpc-1 (statistical). However, as discussed in [193], several systematic uncertainties are not fully accounted for in the reported error budget. The distance calibration of CF-4 galaxies introduces uncertainties of 2–3%, translating to 1.5–2.3 km s-1 Mpc-1 in . The halo mass matching between simulations and observations, which is crucial for their method, contributes an estimated systematic error of 1–2% (0.8–1.5 km s-1 Mpc-1). Additionally, the modeling of pairwise velocities in the nonlinear regime, where theoretical predictions are less certain, introduces uncertainties of 2–3% (1.5–2.3 km s-1 Mpc-1). The choice of the maximum separation scale ( Mpc) and the selection of the velocity threshold add further systematic uncertainties of order 1% (0.8 km s-1 Mpc-1). Cosmic variance, despite mitigation through multiple simulation realizations, remains a source of systematic error. Combining these contributions in quadrature yields a total systematic uncertainty of approximately 3–4 km s-1 Mpc-1, which is comparable to or larger than the quoted statistical uncertainty. Therefore, a more complete error budget would be km s-1 Mpc-1, or equivalently km s-1 Mpc-1 when statistical and systematic errors are combined in quadrature. This corrected level of uncertainty is used in our analysis.
B.2 Key DLI–SHF Measurements
@@ B.2.1 TDCOSMO: Evolution of Strong Lensing Constraints
TDCOSMO IV (Birrer et al. 2020): [21] Incorporating stellar kinematics to constrain the mass-sheet degeneracy (MSD) shifted the result to km s-1 Mpc-1—now consistent with Planck. This dramatic shift demonstrated that power-law profiles alone cannot break the MSD.
Recent results (2023–2025): Subsequent analyses using hierarchical modeling, additional lens systems, and improved kinematics yield in the range – km s-1 Mpc-1, depending on modeling assumptions [163, 175].
@@
We exclude the original Pesce et al. measurement from our sample, as the more recent Barua et al. analysis uses the same underlying data while addressing the known systematic uncertainties of the earlier measurement.
B.2.2 Liu et al. (2023): SDSS J1004+4112 Cluster Lens
The Liu et al. [104] analysis of the cluster-lensed quasar SDSS J1004+4112 explores 16 different mass models, yielding values spanning 55–90 km s-1 Mpc-1. Their combined constraint with equal weighting gives km s-1 Mpc-1.
In our analysis, we adopt their more constrained result of km s-1 Mpc-1, which corresponds to their selection of the two mass models (m113 and m123) that best reproduce the observed shapes of the lensed quasar host galaxy and background galaxies (see their Figure 5 and Section V). The authors themselves emphasize this as a more physically motivated constraint, stating that these models “best match the observed shapes.”
We acknowledge that this choice favors the lower value. However, selecting mass models based on their ability to reproduce independent observables (lensed galaxy morphologies) is more principled than equal weighting of all models regardless of their physical plausibility. Readers should note that the conservative combined result ( km s-1 Mpc-1) remains consistent with our conclusions given its large uncertainty.
B.2.3 Kelly et al. (2023) and Li et al. (2024): SN Refsdal Time Delays
@@
Supernova Refsdal, the first multiply-imaged supernova [86], provides unique advantages for time-delay cosmography: the supernova light curve allows unambiguous identification of multiple images, unlike quasar variability.
Both Kelly et al. [87] and Li et al. [101] analyze time delays from SN Refsdal, yet we classify them in different categories based on their methodological approaches:
@@
Kelly et al. (2023) — Category 2: This analysis uses time delays from the “Einstein Cross” configuration plus the predicted reappearance of image “SX.” The result, km s-1 Mpc-1, is derived using specific cluster mass models that assume CDM cosmology for the time-delay distance calculation and interprets results within the CDM framework.
Li et al. (2024) — Category 3: This analysis explicitly develops a “cosmological-model-independent” approach, as emphasized in their title and methodology. They obtain km s-1 Mpc-1 using a combination of strong lensing and other probes to determine without assuming CDM.
The distinction hinges on whether the original analysis explicitly states model-independence. While both measurements yield consistent values and analyze the same underlying data, their category assignments follow our classification criterion based on the authors’ stated assumptions. This example illustrates that identical data can yield measurements in different categories depending on the analysis methodology employed.
B.2.4 Piratova-Moreno et al.:FRBs
@@
Piratova-Moreno et al. (2025) [140] employed Fast Radio Bursts (FRBs) as cosmological probes to measure the Hubble constant using a catalog of 98 localized FRBs. Using the Maximum Likelihood Estimate (MLE) method with confirmed FRBs, they obtained km s-1 Mpc-1 (statistical), while their mock FRB catalog yielded km s-1 Mpc-1. However, this method is subject to substantial systematic uncertainties that are not fully quantified in the reported error budget. The primary systematic error arises from the decomposition of the observed dispersion measure (DM) into its constituent components: the Milky Way interstellar medium and halo (), the intergalactic medium (), and the host galaxy (). The uncertainty in depends on the Galactic electron density model used (NE2001 vs. YMW16), with differences of 20% at low Galactic latitudes, contributing 1.5–2 km s-1 Mpc-1 to the uncertainty. The host galaxy contribution, , is highly uncertain and varies dramatically between FRBs—from 30 pc cm-3 for typical hosts to over 900 pc cm-3 for extreme cases like FRB 20190520B. Studies suggest uncertainties of 50–100 pc cm-3, which translate to systematic errors of 2–3 km s-1 Mpc-1 in . Additionally, the baryon fraction in the IGM () introduces further systematic uncertainty; recent estimates range from to , corresponding to 1–2 km s-1 Mpc-1 uncertainty in . The choice of DM– relation (linear vs. power-law) also significantly impacts results, with Piratova-Moreno et al. reporting values ranging from km s-1 Mpc-1 (linear) to km s-1 Mpc-1 (power-law), indicating substantial model-dependent systematic uncertainties of order 5–15 km s-1 Mpc-1. Combining these systematic contributions in quadrature yields a total systematic error of approximately 6–8 km s-1 Mpc-1 for the MLE result, dominating over the statistical uncertainty. A more realistic error budget would therefore be km s-1 Mpc-1, or equivalently km s-1 Mpc-1 when combined in quadrature, reflecting the current limitations of FRB-based cosmology. This is the value used in our analysis.
B.2.5 Vogl et al. (2024): Tailored EPM
@@
The Vogl et al. [182] tailored Expanding Photosphere Method represents the most precise cosmological-model-independent measurement supporting the high- solution outside the traditional distance ladder.
@@
B.2.6 Barua et al. (2025): Megamaser Cosmology Project
The Megamaser Cosmology Project measures geometric distances to galaxies with circumnuclear water maser disks. The Barua et al. [17] frequentist analysis yields km s-1 Mpc-1.
@@
B.2.7 Pogosian et al. (2024): Sound-Horizon-Free BAO
The Pogosian et al. [144] analysis demonstrates that can be constrained from BAO data without assuming the CMB Sound Horizon.
@@
Results: DESI BAO DR2 yields ; pre-DESI BAO yields . The DESI DR2 result was presented by L. Pogosian during the CosmoVerse Seminar Series [145] and represents an update of [144] using DESI DR1 BAO data; publication is forthcoming.
Significance: These measurements anchor the low- solution in Category 4. Their high precision (%) means they dominate the weighted mean of this category.
B.2.8 Gravitational Wave standard sirens
GW standard sirens provide constraints independent of electromagnetic distance ladder:
Bright sirens (GW170817): The binary neutron star merger with optical counterpart yields km s-1 Mpc-1 [122], limited by the peculiar velocity of NGC 4993 and viewing angle degeneracy. Additionally, Wang et al. (2025) [185] find a consistent value of km s-1 Mpc-1. These measurements provide important complementary constraints that are independent of the Distance Ladder.
Dark sirens (GWTC-3): Statistical analysis of 47 events without counterparts yields km s-1 Mpc-1 [102], consistent with Planck.
@@ B.2.8 Gravitational Wave standard sirens
B.2.9 Fast Radio Burst Measurements
@@
Galactic electron density models
Results: Published values span – km s-1 Mpc-1. Maximum-likelihood approaches [140] tend to favor lower values ( km s-1 Mpc-1). Also the estimation of using Fast Radio Bursts (FRBs) is highly sensitive to the model adopted for the Galactic electron density distribution. As highlighted by recent studies (e.g., Wang et al. [184]), switching between established models (such as NE2001 and YMW16) can shift the inferred value significantly (e.g., from to km s-1 Mpc-1), introducing a systematic uncertainty that rivals the statistical error.As FRB samples grow and host DM distributions become better characterized, this probe may provide competitive constraints.
@@ C.1.1 Distance Ladder Correlations
SNe Ia samples: Hubble-flow SNe Ia are predominantly from Pantheon+ [25], with substantial overlap. standardization parameters (, , ) are often shared.
@@ C.2 Methodology for Correlation-Adjusted Estimates
where is the covariance matrix with elements . For simplicity, we adopt a uniform correlation structure within each category:
@@ C.3.1 Distance Ladder (Category 1)
compared to km s-1 Mpc-1 under independence. The central value shifts by only km s-1 Mpc-1 (within statistical uncertainty), but the uncertainty increases by %.
C.3.2 Combined (Categories 2+3+4)
@@
compared to km s-1 Mpc-1 under independence. The central value remains essentially unchanged, and the uncertainty increases by %.